# 22footnotetext: Corresponding author

### v1

@@ 5.1. Segmentation
Table 9. Quantitative Comparison of Various Segmentation Models across Different Datasets
@@
SegMamba (Xing et al., 2024) BraTS2023 (Menze et al., 2015; et al., 2024b; Bakas et al., 2017) 3.56 - 91.32 - - - - - - - - -
AIIB202 (Nan et al., 2022; Li et al., 2022) - 88.59 - - - - - - - - - -
CRC-500 dataset (Xing et al., 2024) 30.89 - 48.02 - - - - - - - - -
H-vmunet(VSS) (Wu et al., 2024a) ISIC2017 (Codella et al., 2018) - - 90.68 88.97 98.23 96.42 - - - - - -
Spleen (et al., 2022, 2019a) - - 94.03 93.30 99.92 99.82 - - - - - -
CVC-ClinicDB (Bernal et al., 2015) - - 89.84 87.68 99.21 98.13 - - - - - -
H-vmunet(H-VSS) (Wu et al., 2024a) ISIC2017 (Codella et al., 2018) - - 91.72 90.56 98.31 96.80 - - - - - -
Spleen (et al., 2022, 2019a) - - 95.71 96.42 99.92 99.87 - - - - - -
CVC-ClinicDB (Bernal et al., 2015) - - 90.87 88.03 99.40 98.33 - - - - - -
UltraLight VM-UNet (Wu et al., 2024b) ISIC2017 (Codella et al., 2018) - - 90.91 90.53 97.90 96.46 - - - - - -
ISIC2018 (Codella et al., 2019; Tschandl et al., 2018) - - 89.40 86.80 97.81 95.58 - - - - - -
${PH}^{\mathbf{2}}$ Dataset (Mendonça et al., 2013) - - 92.65 93.45 96.06 95.21 - - - - - -
HC-Mamba (Xu, 2024) ISIC2017 (Codella et al., 2018) - - 88.18 95.17 97.47 86.99 79.27 - - - - -
ISIC2018 (Codella et al., 2019; Tschandl et al., 2018) - - 89.25 87.90 97.08 94.84 - - - - -
@@
U-Mamba_Bot (Ma et al., 2024) 3D Abdomen CT dataset (et al., 2023c) - - 86.83±08.08 - - - - 90.49±08.21 - - - -
2D Abdomen MRI dataset (Ji et al., 2022) - - 75.88±10.51 - - - - 82.85±10.74 - - - -
Endoscopy dataset (et al., 2019b) - - 65.40±30.08 - - - - 66.92±30.50 - - - -
Microscopy dataset (Ma et al., 2024) - - - - - - - - 53.89±28.17 - - -
U-Mamba_Enc (Ma et al., 2024) 3D Abdomen CT dataset (et al., 2023c) - - 86.38±09.08 - - - - 89.80±09.21 - - - -
2D Abdomen MRI dataset (Ji et al., 2022) - - 76.25±10.82 - - - - 83.27±10.87 - - - -
Endoscopy dataset (et al., 2019b) - - 63.03±30.67 - - - - 64.51±31.04 - - - -
Microscopy dataset (Ma et al., 2024) - - - - - - - - 56.07±27.84 - - -
LKM-UNet (Wang et al., 2024a) 3D Abdomen CT dataset (et al., 2023c) - - 86.82 - - - - 90.02 - - - -
2D Abdomen MRI dataset (Ji et al., 2022) - - 77.35 - - - - 83.80 - - - -
Swin-UMamba (Liu et al., 2024b) 2D Abdomen MRI Dataset (Ji et al., 2022) - - 77.60 - - - - 84.21 - - - -
Endoscopy Dataset (et al., 2019b) - - 67.67 - - - - 69.22 - - - -
Microscopy Dataset (Ma et al., 2024) - - - - - - - - 58.06 - - -
Swin-UMamba (Liu et al., 2024b) 2D Abdomen MRI Dataset (Ji et al., 2022) - - 77.05 - - - 83.76 - - - -
(Using Decoder) Endoscopy Dataset (et al., 2019b) - - 67.83 - - - - 69.33 - - - -
Microscopy Dataset (Ma et al., 2024) - - - - - - - - 59.82 - - -
LightM-UNet (Liao et al., 2024) LiTS Dataset (et al., 2023d) - - 84.58 - - - 77.48 - - - - -
Montgomery&Shenzhen Dataset (Jaeger et al., 2014) - - 96.17 - - - 92.74 - - - - -
UU-Mamba (Tsai et al., 2024) ACDC MRI Cardiac Dataset(Bernard et al., 2018) - - 92.79 - - - - - - - - -
Mamba-HUNet (Sanjid et al., 2024) Brain MRI Multiple 2.25 85.36 92.87 92.94 98.65 - - - - - - -
Sclerosis Dataset (M. Muslim e al., 2022)
Vm-Unet (Ruan and Xiang, 2024) Brain MRI Multiple 2.9 80.39 87.49 91.35 95.31 - - - - - - -
Sclerosis Dataset (M. Muslim e al., 2022)
Mamba-UNet (Wang et al., 2024d) ACDC MRI Cardiac Dataset(Bernard et al., 2018) - 86.98 92.81 92.89 98.59 99.72 - - - 2.46 0.76 92.75
@@
Weak-Mamba-Unet (Wang and Ma, 2024) ACDC MRI Cardiac Dataset(Bernard et al., 2018) - - 91.71 93.09 99.20 99.63 - - - 3.95 0.88 90.95
Semi-Mamba-UNet (Ma and Wang, 2024) ACDC MRI Cardiac Dataset(Bernard et al., 2018)
@@
CAMS-Net (Khan et al., 2024) MICCAI-2023 (et al., 2024c) - - 84.84 - - - - - - 6.51 - -
MUCM-Net (Yuan et al., 2024) ISIC2017 (Codella et al., 2018) - - 91.85 90.14 98.57 96.97 - - - - - -
ISIC2018 (Codella et al., 2019; Tschandl et al., 2018) - - 90.95 90.46 97.72 96.18 - - - - - -
@@
5.2. Classification
In this section, we present Table 10 which shows a detailed comparison of Mamba-based models that perform classification task across different datasets. The metrics containes Accuracy (Acc), F1-score and Area Under the Curve (AUC), where higher values indicate better performance. This table helps to assess the effectiveness of each Mamba models in achieving accurate and reliable classification results in medical imaging.
Table 10. Quantitative Comparison of Mamba-based Models on Classification Tasks
@@ 5.2. Classification
Microscopic-Mamba RPE Data (et al., 2016) - - 98.17
MHIST (Wei et al., 2021) - - 94.17
SARS (H et al., 2023) - - 99.48
Tissue MNIST (Yang et al., 2023) - - 93.50
Med FM Colon (Wang et al., 2023) - - 99.64
@@ 5.3. Registration
Table 11. Quantitative Comparison of Mamba-based Image Registration Models on Testing set of SR-Reg (SynthRAD Registration) (Guo et al., 2024) Dataset
Methods Dice(%) $\uparrow$ HD95(mm) $\downarrow$ $P|J\phi|\leq$ 0(%) $\downarrow$ Time Memory Parameter
@@ 6.1. Limitations
6.2. Emerging Areas
Mamba 2:
Figure 30. (a) Relationship between SSMs and Attention through Structured Matrices , (b) Structured State Space Duality (Dao and Gu, 2024)
@@ 6.2. Emerging Areas
Mamba 2 proposed by Dao et al. (Dao and Gu, 2024) reduces the gap between recurrent nature of SSM and parallel nature of attention in transformers. Figure 30 shows Structured State Space Duality which explains the relationship between SSM and attention using structures matrices to describe their connection. The recurrent state space equations are converted into matrix form by expanding recurrence equations commonly used in SSM. This approach resembles the kernel representation utilized in Gu et al. (Gu et al., 2021a). The conversion of the recurrent equations of Gu et al. (Gu and Dao, 2024) involves matrix representations that have been discretized using the Zero Order Hold (ZOH) method, as previously discussed in this work. Both the linear (recurrent) and quadratic (attention) forms are unified under the framework known as State Space Dual (SSD) Layer. In the recurrent form, the parameter $A$ from Mamba is simplified from a diagonal to a scalar times identity structure, while larger head dimensions $P$ are employed for this model. The authors introduce a 1SS (a) mask, derived from unrolling recurrence equations of SSM, which simplifies the parameter $A$ when attention equations are rearranged into linear attention forms. This reduced attention equation is termed Structured Masked Attention (SMA). Using SMA, the authors propose architectures equivalent to Multi-Head Attention (MHA), termed Multi-Head SSM, and Multi-Query Attention, known as Multi-Contact SSM. Additionally, Multi-Contact SSM can be extended to include Group Query Attention.
Figure 31. Comparison of Mamba and Mamba-2 Architecture (Dao and Gu, 2024)
With the established relationship between transformers and SSM, models now leverage parallelism and equivalent implementations similar to ViT. This approach enables the creation of more expressive models using minimal parameters. Figure 31 compares the sequential block of Mamba and parallel block of Mamba-2 architecture.
@@
xLSTM: Extended Long Short-Term Memory (xLSTM) proposed by Beck et al. (Beck et al., 2024) incorporates sLSTM and mLSTM, sLSTM have proposed a new normalized gating mechanism compared to original LSTM and also changed gating mechanism from sigmoid to exponential gating. mLSTM has a learnable projection matrix for query, key, value similar to transformer (Vaswani et al., 2017), and transforms scalar memory states into matrix-based representations for parallel training. Vision-xLSTM proposed by Alkin et al. (Alkin et al., 2024) employs stacked mLSTM blocks similar to ViT, demonstrating superior performance over its Mamba counterpart (ViM) while maintaining minimal parameter complexity. xLSTM-Unet proposed by Chen et al. (Chen et al., 2024b) utilizes a U-Net-based model for segmenting 2D MRI, endoscopy, and microscopy images, outperforming both Mamba-based and Transformer-based models.
@@ 7.2. Summary of Key Findings
7.3. Future Directions
Figure 32. Future Directions of Mamba
@@ 7.3. Future Directions
Mamba’s computational efficiency, similar to that of CNNs, enables it to perform well even without large-scale datasets, making it a strong candidate for downstream tasks, multi-tasks, and pre-trained model adaptations. Figure 32 describes future directions of Mamba and Mamba-based architectures in medical applications. Its design, including the use of SSMs, reduces computational complexity, making mamba particularly suited for processing high-resolution data such as remote sensing images, whole slide images, and extended video sequences. In multimodal settings, however, Mamba faces challenges in uniformly learning features across different data types. Similar to transformers which handle both text and images, Mamba’s capacity for processing extended sequences offers potential for multimodal learning. To fully leverage this potential, an efficient approach is needed, and Mamba2 aims to address these issues. In-context learning has also become more sophisticated, and Mamba’s ability to model long-range dependencies makes it promising for improving performance across NLP, CV, and multimodal domains. Mamba’s selective scanning mechanism requires adaptation for non-causal characteristics of visual data. New scanning techniques are necessary to fully exploit higher-dimensional non-causal visual data. In Mamba 2, the scanning process is made parallelly causal using Semiseparable SMA, which models autoregressively in a way similar to decoder-only transformers such as GPT (Tan et al., 2021; et al., 2020) and Llama (Touvron et al., 2023; et al., 2023b, 2024a).

### v3

@@ 5.1 Segmentation
Table 9: Quantitative Comparison of Various Segmentation Models across Different Datasets
@@
SegMamba [16] BraTS2023 [41, 42, 43] 3.56 - 91.32 - - - - - - - - -
AIIB202 [44, 45] - 88.59 - - - - - - - - - -
CRC-500 dataset [16] 30.89 - 48.02 - - - - - - - - -
H-vmunet(VSS) [17] ISIC2017 [46] - - 90.68 88.97 98.23 96.42 - - - - - -
Spleen [164, 165] - - 94.03 93.30 99.92 99.82 - - - - - -
CVC-ClinicDB [49] - - 89.84 87.68 99.21 98.13 - - - - - -
H-vmunet(H-VSS) [17] ISIC2017 [46] - - 91.72 90.56 98.31 96.80 - - - - - -
Spleen [164, 165] - - 95.71 96.42 99.92 99.87 - - - - - -
CVC-ClinicDB [49] - - 90.87 88.03 99.40 98.33 - - - - - -
UltraLight VM-UNet [81] ISIC2017 [46] - - 90.91 90.53 97.90 96.46 - - - - - -
ISIC2018 [47, 48] - - 89.40 86.80 97.81 95.58 - - - - - -
${PH}^{\mathbf{2}}$ Dataset [166] - - 92.65 93.45 96.06 95.21 - - - - - -
HC-Mamba [83] ISIC2017 [46] - - 88.18 95.17 97.47 86.99 79.27 - - - - -
ISIC2018 [47, 48] - - 89.25 87.90 97.08 94.84 - - - - -
@@
U-Mamba_Bot [86] 3D Abdomen CT dataset [50] - - 86.83±08.08 - - - - 90.49±08.21 - - - -
2D Abdomen MRI dataset [51] - - 75.88±10.51 - - - - 82.85±10.74 - - - -
Endoscopy dataset [52] - - 65.40±30.08 - - - - 66.92±30.50 - - - -
Microscopy dataset [53] - - - - - - - - 53.89±28.17 - - -
U-Mamba_Enc [86] 3D Abdomen CT dataset [50] - - 86.38±09.08 - - - - 89.80±09.21 - - - -
2D Abdomen MRI dataset [51] - - 76.25±10.82 - - - - 83.27±10.87 - - - -
Endoscopy dataset [52] - - 63.03±30.67 - - - - 64.51±31.04 - - - -
Microscopy dataset [53] - - - - - - - - 56.07±27.84 - - -
LKM-UNet [18] 3D Abdomen CT dataset [50] - - 86.82 - - - - 90.02 - - - -
2D Abdomen MRI dataset [51] - - 77.35 - - - - 83.80 - - - -
Swin-UMamba [110] 2D Abdomen MRI Dataset [51] - - 77.60 - - - - 84.21 - - - -
Endoscopy Dataset [52] - - 67.67 - - - - 69.22 - - - -
Microscopy Dataset [53] - - - - - - - - 58.06 - - -
Swin-UMamba [110] 2D Abdomen MRI Dataset [51] - - 77.05 - - - 83.76 - - - -
(Using Decoder) Endoscopy Dataset [52] - - 67.83 - - - - 69.33 - - - -
Microscopy Dataset [53] - - - - - - - - 59.82 - - -
LightM-UNet [80] LiTS Dataset [55] - - 84.58 - - - 77.48 - - - - -
Montgomery&Shenzhen Dataset [56] - - 96.17 - - - 92.74 - - - - -
UU-Mamba [88] ACDC MRI Cardiac Dataset[54] - - 92.79 - - - - - - - - -
Mamba-HUNet [87] Brain MRI Multiple 2.25 85.36 92.87 92.94 98.65 - - - - - - -
Sclerosis Dataset [57]
Vm-Unet [115] Brain MRI Multiple 2.9 80.39 87.49 91.35 95.31 - - - - - - -
Sclerosis Dataset [57]
Mamba-UNet [78] ACDC MRI Cardiac Dataset[54] - 86.98 92.81 92.89 98.59 99.72 - - - 2.46 0.76 92.75
@@
Weak-Mamba-Unet [79] ACDC MRI Cardiac Dataset[54] - - 91.71 93.09 99.20 99.63 - - - 3.95 0.88 90.95
Semi-Mamba-UNet [84] ACDC MRI Cardiac Dataset[54]
@@
CAMS-Net [89] MICCAI-2023 [163] - - 84.84 - - - - - - 6.51 - -
MUCM-Net [90] ISIC2017 [46] - - 91.85 90.14 98.57 96.97 - - - - - -
ISIC2018 [47, 48] - - 90.95 90.46 97.72 96.18 - - - - - -
@@
5.2 Classification
In this section, we present Table 10 which shows a detailed comparison of Mamba-based models that perform classification task across different datasets. The metrics containes Accuracy (Acc), F1-score and Area Under the Curve (AUC), where higher values indicate better performance [167, 168]. This table helps to assess the effectiveness of each Mamba models in achieving accurate and reliable classification results in medical imaging.
Table 10: Quantitative Comparison of Mamba-based Models on Classification Tasks
@@ 5.2 Classification
Microscopic-Mamba RPE Data [169] - - 98.17
MHIST [170] - - 94.17
SARS [171] - - 99.48
Tissue MNIST [61] - - 93.50
Med FM Colon [172] - - 99.64
@@ 5.3 Registration
Table 11: Quantitative Comparison of Mamba-based Image Registration Models on Testing set of SR-Reg (SynthRAD Registration) [66] Dataset
Methods Dice(%) $\uparrow$ HD95($m$) $\downarrow$ $P|J\phi|\leq$ 0(%) $\downarrow$ Time Memory Parameter
@@ 6.1 Limitations
  5. 5.
    Multimodal Model Complexity and Resource Demands: Multimodal Mamba-based models use several specialized components, like Learnable Descriptive Convolution and Cross Modality Fusion Mamba, along with multi-step processes such as combining 3D GANs with Mamba classifiers. This makes the models quite complex, which can make them harder to build, train, and maintain. Even though they avoid heavy self-attention, these models still have many deep layers and fusion operations that require a lot of computing power. As a result, they can be slow and less practical for real-time or edge applications in clinical settings.
  6. 6.
    Self-Supervised Learning Challenges: Self-supervised Mamba models also tend to be complicated and require significant computational resources. They need careful setup and large amounts of unlabeled data to train effectively. Because of the long training times and resource demands, using these models in many real-world medical environments can be challenging.
6.2 Emerging Areas
The emerging areas includes Mamba 2 and xLSTM. Due to the recurrent nature of Mamba, the classical RNN has also evolved resulting in variants such as Min RNN, Min LSTM and Min GRU which enable these algorithms to perform parallel scan [173].
@@ 6.2 Emerging Areas
Mamba 2: Mamba 2 proposed by Dao et al. [39] reduces the gap between recurrent nature of SSM and parallel nature of attention in transformers. Figure 18 shows Structured State Space Duality which explains the relationship between SSM and attention using structures matrices to describe their connection. The recurrent state space equations are converted into matrix form by expanding recurrence equations commonly used in SSM. This approach resembles the kernel representation utilized in Gu et al. [14]. The conversion of the recurrent equations of Gu et al. [15] involves matrix representations that have been discretized using the Zero Order Hold (ZOH) method, as previously discussed in this work. Both the linear (recurrent) and quadratic (attention) forms are unified under the framework known as State Space Dual (SSD) Layer. In the recurrent form, the parameter $A$ from Mamba is simplified from a diagonal to a scalar times identity structure, while larger head dimensions $P$ are employed for this model. The authors introduce a 1SS (a) mask, derived from unrolling recurrence equations of SSM, which simplifies the parameter $A$ when attention equations are rearranged into linear attention forms. This reduced attention equation is termed Structured Masked Attention (SMA). Using SMA, the authors propose architectures equivalent to Multi-Head Attention (MHA), termed Multi-Head SSM, and Multi-Query Attention, known as Multi-Contact SSM. Additionally, Multi-Contact SSM can be extended to include Group Query Attention. With the established relationship between transformers and SSM, models now leverage parallelism and equivalent implementations similar to ViT. This approach enables the creation of more expressive models using minimal parameters.
Figure 18: (a) Relationship between SSMs and Attention through Structured Matrices , (b) Structured State Space Duality [39]
@@
xLSTM: Extended Long Short-Term Memory (xLSTM) proposed by Beck et al. [40] incorporates sLSTM and mLSTM, sLSTM have proposed a new normalized gating mechanism compared to original LSTM and also changed gating mechanism from sigmoid to exponential gating. mLSTM has a learnable projection matrix for query, key, value similar to transformer [12], and transforms scalar memory states into matrix-based representations for parallel training. Vision-xLSTM proposed by Alkin et al. [174] employs stacked mLSTM blocks similar to ViT, demonstrating superior performance over its Mamba counterpart (ViM) while maintaining minimal parameter complexity. xLSTM-Unet proposed by Chen et al. [175] utilizes a U-Net-based model for segmenting 2D MRI, endoscopy, and microscopy images, outperforming both Mamba-based and Transformer-based models.
@@ 7.2 Summary of Key Findings
Variants of U-Net incorporated Mamba blocks to improve medical image segmentation by integrating these blocks at different stages of architecture such as before encoder layers, within skip connections or replacing encoder blocks. This improves feature extraction while retaining U-Net’s core strengths. These designs boost both global and local feature modeling by capturing long-range dependencies and spatial details through bidirectional scanning and large receptive fields. Some variants replace traditional skip connections with fusion modules that combines low-level and high-level features to understand complex image patterns. Overall, these U-Net variants achieve better segmentation accuracy and efficiency which makes them suitable for large-scale medical imaging tasks.
A key strength of Mamba lies in its scanning methods, which work well for medical data with spatial and temporal dimensions. Unlike self-attention, which has quadratic time complexity, scan-based operations scale linearly, making them more efficient for high-resolution inputs. Mamba uses various techniques such as bidirectional, selective, continuous 2D, zigzag, spatiotemporal, local and efficient 2D scans to capture patterns across directions and modalities while preserving spatial continuity. Extensions like multi-path, hierarchical, multi-head and omnidirectional scans further improve feature extraction by enabling diverse spatial and temporal interactions. These strategies reduce computational cost and improve the model’s ability to interpret complex medical images and videos with greater detail. The review also shows that Mamba architectures play a key role in improving medical image analysis, especially when labeled data is scarce or imperfect. They are strong at capturing complex spatial relationships and long-range patterns, which helps them learn effectively from noisy or limited labels. Techniques like pixel-level contrastive learning and cross-supervised training boost the model’s ability to extract meaningful features, leading to more accurate segmentation and classification results. Overall, Mamba-based methods offer a practical solution to common challenges in medical imaging, reducing the need for large annotated datasets while enhancing diagnostic performance.
Additionally, multimodal Mamba models integrate various types of data such as images, lab tests and genetic information. This makes them especially useful for capturing richer patterns that single-modality models often miss. Their structure is also better suited for learning long-range relationships across modalities compared to older CNN or Transformer-based approaches, which can struggle with efficiency or complexity. The ability of Mamba to handle complex medical tasks using diverse inputs can help improve the accuracy of diagnosis and support more personalized treatment. Multimodal architectures also show promise for adapting to a wide range of clinical applications and evolving medical data types.
7.3 Future Directions
@@
Mamba’s computational efficiency, similar to that of CNNs, enables it to perform well even without large-scale datasets, making it a strong candidate for downstream tasks, multi-tasks, and pre-trained model adaptations. Its design, including the use of SSMs, reduces computational complexity, making mamba particularly suited for processing high-resolution data such as remote sensing images, whole slide images, and extended video sequences. In multimodal settings, however, Mamba faces challenges in uniformly learning features across different data types. Similar to transformers which handle both text and images, Mamba’s capacity for processing extended sequences offers potential for multimodal learning. To fully leverage this potential, an efficient approach is needed, and Mamba2 aims to address these issues. In-context learning has also become more sophisticated, and Mamba’s ability to model long-range dependencies makes it promising for improving performance across NLP, CV, and multimodal domains. Mamba’s selective scanning mechanism requires adaptation for non-causal characteristics of visual data. New scanning techniques are necessary to fully exploit higher-dimensional non-causal visual data. In Mamba 2, the scanning process is made parallelly causal using Semiseparable SMA, which models autoregressively in a way similar to decoder-only transformers such as GPT [176, 177] and Llama [178, 179, 180].